OpenStudy (anonymous):

What is the fourth term in this sequence? tn=3.2n-5 i dont get it...

OpenStudy (tkhunny):

\(t_{0} = 3.2(0) - 5 = -5\) \(t_{1} = 3.2(1) - 5 = -1.8\) \(t_{2} = 3.2(2) - 5 = 1.4\) etc. It is possible that you are to start with n = 1, rather than n = 0 as I demonstrated, above.

OpenStudy (anonymous):

ok thanks:)

OpenStudy (anonymous):

@tkhunny the answers are -7.2 -16.8 7.8 though

OpenStudy (anonymous):

@Frostbite @AriPotta

OpenStudy (tkhunny):

Keep going: \(t_{3} =3.2(3)−5=4.6\) \(t_{4} =3.2(4)−5=7.8\) And I think we're done.

OpenStudy (aripotta):

you have to substitute 4 for n

OpenStudy (anonymous):

oooooh thanks:)